# T-3DGS: Removing Transient Objects for 3D Scene Reconstruction

### v1

@@ 4 Experiments
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
NeRF On-the-go [31] 20.15 0.64 20.11 0.61 24.22 0.81 20.78 0.75 23.33 0.79 21.41 0.72 21.67 0.72
3DGS [12] 20.09 0.70 20.11 0.69 23.51 0.85 15.87 0.68 20.78 0.80 17.80 0.69 19.69 0.74
Robust3DGS [43] 16.97 0.61 18.18 0.59 23.47 0.85 21.33 0.85 22.61 0.88 21.81 0.82 20.73 0.77
WildGaussians [15] 20.77 0.70 20.74 0.67 25.79 0.88 21.77 0.85 24.39 0.88 22.36 0.80 22.64 0.80
SpotLessSplats [34] 21.25 0.66 20.49 0.63 25.59 0.85 21.13 0.80 24.13 0.78 22.18 0.76 22.46 0.75
Ours 20.89 0.72 21.01 0.69 26.64 0.90 21.55 0.86 25.96 0.91 23.18 0.84 23.21 0.82
Table 1: Quantitative comparison on On-the-go dataset [31].
@@
3DGS [12] 24.49 0.91 20.42 0.87 20.08 0.89 20.45 0.86 26.96 0.94 22.48 0.89
Robust3DGS [43] 25.35 0.93 24.74 0.92 24.33 0.93 22.95 0.91 28.52 0.96 25.18 0.93
WildGaussians [15] 25.71 0.92 23.68 0.91 24.65 0.92 21.69 0.89 28.89 0.95 24.92 0.92
SpotLessSplats [34] 25.28 0.91 24.63 0.90 24.11 0.91 22.22 0.90 28.08 0.92 24.86 0.91
Ours w/o TMR 25.69 0.93 24.74 0.92 23.82 0.92 24.21 0.92 29.01 0.95 25.49 0.93
Ours w/ TMR 27.71 0.95 25.42 0.93 28.34 0.97 28.79 0.97 29.87 0.96 28.03 0.96
Table 2: Quantitative comparison on our T-3DGS dataset.
@@
We evaluate our proposed T-3DGS model on various datasets captured in uncontrolled settings and filled with diverse distractors. We perform qualitative and quantitative comparisons of our method with other SOTA methods. Finally, we provide an ablation study of our design and loss choices and discuss the limitations of the proposed method.
@@
Datasets. We evaluate our model on three challenging datasets. The NeRF On-the-go dataset [31] contains 4 outdoor and 2 indoor sparsely captured scenes with different levels of occlusions (from 5% to over 30%) and minimal changes in appearance. The RobustNeRF dataset [33] contains 5 indoor scenes with unintentional changes during the capture process. Those changes include both transient objects that appear and disappear without forming a coherent video sequence (i.e., do not have a clear temporal ordering) and dynamic objects (e.g., floating balloons). Additionally, we introduce our novel T-3DGS dataset. The dataset contains 5 densely captured indoor scenes. Generally, dynamic objects in our videos are walking people and various small objects. However, unlike previous datasets, all scenes include various adversarial cases, such as transient, semi-transient, and slow-moving objects.
@@
Baselines. We compare our model against vanilla 3D Gaussian Splatting [12] and against the current state-of-the-art method SpotLessSplats [34]. We further include WildGaussians [15] and Robust3DGS [43] as baselines. To compare different models we adopt commonly used PSNR and SSIM [45] metrics for evaluation.
@@
Implementation details. All our experiments are conducted following the training setup from the official 3DGS implementation. We train our models for 30K iterations, using the Adam optimizer with a learning rate of 1e-3 for the TMP. The depth regularization loss $\mathcal{L}_{\text{depth}}$ is activated after the first 500 iterations, allowing the 3DGS to establish initial geometry reconstruction. For the experiments with mask propagation, we first train the TMP for 7000 iterations. At that point we pause the training to propagate the transient masks. Subsequently, we initiate a new training procedure using the propagated masks, keeping all other parameters the same as the original training setup. Thanks to our consistency term, we achieve robust results with a fixed regularization parameter $\lambda_{prior}=0.1$ across all scenes, eliminating the need for per-scene parameter tuning that was required in previous approaches.
Figure 4: Qualitative results on the On-the-go dataset. Our method detects transient objects better than our competitors. Predicted masks are shown in green.
Figure 5: Qualitative results on T-3DGS dataset. Our method detects cleaner transient masks and improves them further by mask propagation TMR module. Predicted masks are shown in green.
@@ 4.1 Quantitative Comparisons
We evaluate our model on all three datasets. We report results on On-the-go and T-3DGS datasets in Tab. 1 and 2, respectively, and we move evaluation results on RobustNeRF dataset to Supplementary Material as it is the least challenging. As shown in Tab. 1 and 2 our method generally outperforms current SOTA methods, in particular, our method is robust to changes in distant and high-frequency details. In Tab. 1 we run our method directly on masks predicted by TMP module without mask propagator.
While current SOTA methods struggle to detect semi-transient objects (Tab. 2), our proposed transient network TMP achieves higher performance minimizing false predictions. The integration of the SAM-based mask propagation TMR module further enhances our results in scene reconstruction filled with semi-transient objects, providing more accurate and reliable reconstructions.
@@ 4.2 Qualitative Comparisons
For qualitative comparison, we compare our method to SpotLessSplats [34] and Robust3DGS [43]. Fig. 4 and 5 show that our method has very few false negatives and robustly detects transients. For example, on the On-the-go dataset most of the methods struggle with high-frequency details and distant objects, as both generally reconstruct slower than the rest of the scene, and therefore confuse RGB residuals-based and similar methods. However, due to the introduction of our robust loss, such artifacts generally disappear from our dynamic maps. While our masks are less sharp, as we upsample them from low-resolution DINOv2 patches, techniques such as guided upsampling can improve performance of our method, however we leave this topic to following works. It should be noted, that SpotLessSplats [34] use features obtained from higher-resolution images, while we compute our features on a lower resolution, the same resolution used for training 3DGS.
For our T-3DGS dataset, we additionally utilize the SAM-based mask propagation module to propagate object-aware masks for semi-transient objects as shown in Fig. 5. While most methods would benefit from such a technique in principle, our masks have higher quality and fewer wrong detections. Applying mask propagation to other methods risks propagating errors.
@@
w/o dilation 22.65 0.82
w/o $\mathcal{L}_{\text{consist}}$ 16.76 0.69
w/o $\mathcal{L}_{\text{depth}}$ 23.02 0.82
Ours (full) 23.21 0.82
Table 3: We ablate the importance of each component of our method on On-the-go dataset. We report the mean over all scenes.
@@ 4.3 Ablation Study
We show the importance of each component in our T-3DGS model, separately disabling each of them. We report ablation results in Tab. 3 on On-the-go dataset. In the first experiment, we show the importance of the dilation before upsampling per-feature probability masks $P^{f}$. Without dilation the produced masks keep shrinking during the optimization process. This leads to a quality reduction in the renderings. In the second experiment, we turn off consistency loss term $\mathcal{L}_{\text{consist}}$. Our results show that adding this loss term improves the quality of renderings across all datasets. In our last experiment, we ablate the regularization term on depth $\mathcal{L}_{\text{depth}}$. Even though the metrics are only slightly improved, adding this term helps to remove floating artifacts in the reconstructed scene. We include visualizations to support our statements in Supplementary Material.
@@ 5 Conclusion
In this work, we have presented novel T-3DGS method for 3D scene reconstruction using Gaussian Splatting by effectively removing foreground dynamic distractors from input videos. The key to the success are unsupervised classification network to identify transient distractors through novel consistency loss term and the mask propagation approach that improves the boundary and temporal consistency of the predicted transient masks. Our results show that our method achieves significant improvement over the state-of-the-art methods on densely captured videos populated with semi-transient distractors. We believe our method is a valuable contribution towards widespread usage of 3DGS for robust 3D scene reconstruction from real-world videos captured in uncontrolled settings.
[…]
@@
A Additional Implementation Details
For mask filtering and refinement in Sec. 3.3.1, we set $\lambda_{\text{cov}}^{\text{ref}}=0.7$ for initial mask refinement and $\lambda_{\text{cov}}^{\text{val}}=0.7$ for validation during object filtration. For temporal refinement in Sec. 3.3.2, we set the memory size parameter $N_{m}=10$ , which controls the number of frames considered during mask propagation. For the final mask dilation step, we perform $N_{e}=5$ iterations of binary dilation. Also, the mask merging threshold $\lambda_{merge}$ is set to 0.9, and the stability ratio threshold $\lambda_{SR}$ to 0.08 in Sec. 3.3.3.
B Evaluation on RobustNeRF Dataset
We provide evaluation on the RobustNeRF dataset [33]. As shown in Tab. 4 our method generally outperforms 3DGS [12], Robust 3DGS [43], WildGaussians [15], and shows similar performance comparing to SpotLessSplats [34]. We run our method directly on masks predicted by TMP module without mask propagator (TMR). In general, the dataset does not seem to be challenging enough to highlight the differences between different methods.
@@ B Evaluation on RobustNeRF Dataset
Ours 25.04 0.84 22.90 0.87 35.81 0.95 33.36 0.93 32.09 0.92 29.84 0.90
Table 4: Quantitative comparison on RobustNeRF dataset [33].
@@
Figure 6: Qualitative results on On-the-go dataset on the training frames. Our method produces higher-quality renderings without artifacts.
@@
Figure 7: Qualitative results on On-the-go dataset on the testing frames. Our method produces higher-quality renderings without artifacts.
C More Qualitative Comparisons
@@
For qualitative comparison, we compare our method to SpotLessSplats [34] and Robust3DGS [43]. We provide corresponding renderings for the masks shown in the main paper in Sec. 4.2. Fig. 6 and 7 show reconstructions of several scenes from the On-the-go dataset on training and testing frames, respectively. Currently, most methods can produce fairly good reconstructions and avoid significant artifacts, so generally, most methods produce fairly similar results (at least in the absence of semi-transient objects and other adversarial cases). However, it is notable that compared to other residual-based methods, we avoid misclassifying high-frequency details and similar objects. This improvement can be attributed to our consistency loss term.
Figure 8: We ablate the importance of each component of our method on On-the-go dataset on Corner scene. We provide qualitative comparison for masks (row 1), rendered images (row 2), and rendered depths (row 3).
Figure 9: We ablate the importance of each component of our method on On-the-go dataset on Spot scene. We provide qualitative comparison for masks (row 1), rendered images (row 2), and rendered depths (row 3).
D Additional Ablation Study
We provide visualizations that show the importance of each component in our proposed T-3DGS model on On-the-go dataset [31]. We show our ablation visualizations in Fig. 8 and 9.
Figure 10: The comparison of predicted masks for full and reduced scenes.
@@ D Additional Ablation Study
Figure 11: The comparison of predicted masks and scene reconstructions during movement of semi-transient objects for different frames.
E Handling Semi-Transient Objects

### v2

@@ 4 Experiments
Figure 4: Qualitative results on the On-the-go dataset. Our method outperforms existing approaches in detecting transient objects. Predicted transient masks are shown in green.
Figure 5: Qualitative results on the T-3DGS dataset. Our method produces cleaner transient masks and further refines them using the (TMR) module.
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeRF On-the-go [20] 20.15 0.64 0.26 20.11 0.61 0.31 24.22 0.81 0.19 20.78 0.75 0.22 23.33 0.79 0.19 21.41 0.72 0.24 21.67 0.72 0.24
3DGS [9] 19.40 0.66 0.21 19.96 0.66 0.19 20.90 0.71 0.24 17.48 0.70 0.20 20.77 0.69 0.32 17.29 0.60 0.36 19.30 0.67 0.25
Robust3DGS [28] 16.97 0.61 0.31 18.18 0.59 0.32 23.47 0.85 0.10 21.33 0.85 0.07 22.61 0.88 0.12 21.81 0.82 0.17 20.73 0.77 0.19
WildGaussians [12] 20.77 0.70 0.23 20.74 0.67 0.21 25.79 0.88 0.09 21.77 0.85 0.07 24.39 0.88 0.10 22.36 0.80 0.17 22.64 0.80 0.15
SpotLessSplats [22] 21.25 0.66 0.24 20.49 0.63 0.24 25.59 0.85 0.12 21.13 0.80 0.08 24.13 0.78 0.18 22.18 0.76 0.20 22.46 0.75 0.18
Ours 21.11 0.71 0.22 20.94 0.69 0.21 26.46 0.90 0.12 21.95 0.87 0.10 25.78 0.90 0.12 22.76 0.83 0.17 23.17 0.82 0.16
Table 1: Quantitative comparison on the On-the-go dataset [20].
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS [9] 24.49 0.91 - 20.42 0.87 - 20.08 0.89 - 20.45 0.86 - 26.96 0.94 - 22.48 0.89 -
Robust3DGS [28] 25.35 0.93 0.09 24.74 0.92 0.10 24.33 0.93 0.08 22.95 0.91 0.10 28.52 0.96 0.06 25.18 0.93 0.09
WildGaussians [12] 25.71 0.92 0.08 23.68 0.91 0.11 24.65 0.92 0.09 21.69 0.89 0.09 28.89 0.95 0.04 24.92 0.92 0.08
SpotLessSplats [22] 25.28 0.91 0.08 24.63 0.90 0.09 24.11 0.91 0.09 22.22 0.90 0.10 28.08 0.92 0.05 24.86 0.91 0.08
Ours w/o TMR 25.77 0.93 0.09 24.67 0.92 0.10 24.67 0.93 0.08 24.07 0.92 0.09 29.36 0.95 0.05 25.71 0.93 0.08
Ours w/ TMR 27.76 0.96 0.02 25.54 0.93 0.06 28.25 0.97 0.02 29.01 0.96 0.02 29.85 0.96 0.02 28.08 0.96 0.03
Table 2: Quantitative comparison on the T-3DGS dataset.
@@
We evaluate our proposed T-3DGS model on various datasets captured in uncontrolled settings and filled with diverse distractors. We perform qualitative and quantitative comparisons against state-of-the-art methods. Finally, we provide an ablation study of architectural and loss function choices. We discuss the limitations of the proposed method in the Supplementary Material.
@@
Datasets. We evaluate our model on three challenging datasets. The NeRF On-the-go dataset [20] contains four outdoor and two indoor sparsely captured scenes with different levels of occlusion (from 5% to over 30%) and minimal appearance changes. The RobustNeRF dataset [21] contains five indoor scenes with unintentional changes during the capture process. These changes include transient objects that appear and disappear without a consistent temporal order, as well as dynamic objects (e.g., floating balloons). Additionally, we introduce our novel T-3DGS dataset. The dataset contains 5 densely captured indoor scenes. Generally, dynamic objects in our videos are walking people and various small objects. However, unlike previous datasets, all scenes incorporate challenging cases, including transient, semi-transient, and slow-moving objects.
@@
Baselines. We compare our model against vanilla 3D Gaussian Splatting [9] and the current state-of-the-art method, SpotLessSplats [22]. We further include WildGaussians [12] and Robust3DGS [28] as baselines. To compare different models, we use commonly used PSNR, SSIM [30] and LPIPS metrics for evaluation.
@@
Implementation details. All our experiments are conducted in accordance with the training setup from the official 3DGS implementation. We train our models for 30K iterations, using the Adam optimizer with a learning rate of 1e-3 for the RUP. The depth regularization loss $\mathcal{L}_{\text{depth}}$ is activated after the first 500 iterations, allowing the 3DGS to establish initial geometry reconstruction. For the experiments with mask propagation, we first train the RUP for 7000 iterations. At that point, we pause the training to propagate the transient masks. Subsequently, we initiate a new training procedure using the propagated masks, keeping all other parameters the same as the original training setup. We dilate all our masks by 10 pixels, except for the Patio scene, where we use the original mask due to its low resolution.
@@ 4.1 Quantitative Comparisons
We evaluate our model on all three datasets. We report results on On-the-go and T-3DGS datasets in Tab. 1 and 2, respectively, and we move the evaluation results of RobustNeRF dataset to the Supplementary Material as it presents the least challenge. As shown in Tab. 1 and 2, our method generally outperforms current SOTA methods. In particular, our method is robust to changes in distant and high-frequency details. In Tab. 1 we run our method directly on masks predicted by RUP module without mask propagator.
While current SOTA methods struggle to detect semi-transient objects (Tab. 2), our proposed transient network RUP achieves higher performance by minimizing false predictions. The integration of the SAM-based mask propagation TMR module further enhances our results in scenes containing semi-transient objects, providing more accurate and reliable reconstructions.
@@ 4.2 Qualitative Comparisons
For qualitative evaluation, we compare our method to SpotLessSplats [22], Robust3DGS [28], and WildGaussians [12]. Fig. 4 and 5 demonstrate that our method minimizes false negatives and effectively detects transients. For example, in the On-the-go dataset, most methods struggle with high-frequency details and distant objects, as these elements are typically reconstructed more slowly than the rest of the scene, leading to inaccuracies in RGB residual-based approaches. However, due to our robust loss function, such artifacts are largely eliminated from our dynamic maps. Notably, SpotLessSplats uses features obtained from higher-resolution images, while we extract features at a lower resolution, the same resolution used for training 3DGS.
For our T-3DGS dataset, we additionally utilize the SAM-based mask propagation module to propagate object-aware masks for semi-transient objects, as shown in Fig. 5. Although most methods would theoretically benefit from this technique, our masks are of higher quality and result in fewer incorrect detections. Applying mask propagation to other methods may introduce error propagation, as demonstrated in the Supplementary Material.
@@
GT masks w/o $\mathcal{L}_{\text{depth}}$ 22.84 0.82
GT masks w/ $\mathcal{L}_{\text{depth}}$ and dilation 23.43 0.81
Ours w/o dilation and $\mathcal{L}_{\text{depth}}$ 22.60 0.80
Ours w/o dilation 22.88 0.80
Ours (full) 23.41 0.81
Table 3: We evaluate the importance of each component of our method on the On-the-go dataset. We report the average performance across all scenes.
@@ 4.3 Ablation Study
We present ablation results in Table 3 for the On-the-go dataset, excluding the Patio scene due to its low resolution. We evaluate our method under the following conditions: (1) without mask dilation, (2) without mask dilation and $\mathcal{L}_{\text{depth}}$, and (3) with both components enabled. Additionally, we report results obtained with ground truth masks while separately disabling $\mathcal{L}_{\text{depth}}$ and mask dilation by 10 pixels. Even when using ground truth masks, dilation noticeably enhances performance. This contradicts the assumptions made by NeRF-HuGS [2] and Robust3DGS [28], as exact masks do not yield optimal performance metrics. Furthermore, mask dilation aids RUP training by ensuring that all transient objects are fully covered. We also note that our results align very closely with those obtained using GT masks, suggesting that more challenging datasets are required.
@@ 5 Conclusion
In this work, we have presented the novel T-3DGS method for 3D scene reconstruction using Gaussian Splatting by effectively filtering out foreground dynamic distractors from input videos. By integrating an unsupervised classification network with bivariate uncertainty modeling, KL divergence regularization, and a mask propagation strategy, our method achieves superior temporal coherence and boundary accuracy. Evaluations on both sparsely and densely captured datasets confirm significant improvements over state-of-the-art approaches. We believe our method represents a significant step toward the broader adoption of 3DGS for robust 3D scene reconstruction from real-world videos captured in uncontrolled settings.
@@
Appendix A Limitations
We use features upscaled by FeatUP [6] to compute cosine distance, and while it is better than simple bilinear interpolation, it is relatively slow and gives fairly noisy results. Utilizing alternative ways to measure per pixel errors might improve both speed and accuracy of the method. Additionally, the temporal refinement process is constrained by a memory window of $N_{m}$ frames, which means that if an object disappears for more than $N_{m}$ frames and then reappears, it will be treated as a new instance with a different label. This can lead to inconsistent tracking and potentially affect the filtering process, especially for semi-transient objects that may temporarily leave the scene. Furthermore, our current filtering approach using global coverage scores may incorrectly filter out valid dynamic objects that undergo significant size changes, such as objects moving towards or away from the camera, or those experiencing perspective changes. We leave it as a future work.
Appendix B Additional Implementation Details
In Sec. 3.3.1, for mask filtering and refinement, we set $\lambda_{\text{cov}}^{\text{ref}}=0.7$ for initial mask refinement and $\lambda_{\text{cov}}^{\text{val}}=0.7$ for validation during object filtration. For temporal refinement in Sec. 3.3.2, we set the memory size parameter $N_{m}=10$, which controls the number of frames considered during mask propagation. For the final mask dilation step, we perform $N_{e}=5$ iterations of binary dilation. In addition, the mask merging threshold $\lambda_{merge}$ is set to 0.9, and the stability ratio threshold $\lambda_{SR}$ to 0.08 in Sec. 3.3.3.
Our model consists of repeating blocks. We fist use bilinear interpolation to increase the resolution of our features by two. We then apply a simple 3 by 3 convolutional layer that also decreases feature size by a factor of two. We then apply layer normalization followed by the GELU non-linearity. We repeat this sequence three times. After that we project our features with 1 by 1 convolution to obtain logits. We use softplus for $\sigma_{1}$, $\sigma_{2}$ and tanh for $\rho$. The normalization layer is crucial for improving the numerical stability that arises due to matrix $\Sigma$ being potentially ill-conditioned.
Appendix C Evaluation on RobustNeRF Dataset
We evaluate our method on the RobustNeRF dataset [21]. As shown in Tab. 4 our method generally outperforms 3DGS [9], Robust 3DGS [28], WildGaussians [12], and shows similar performance compared to SpotLessSplats [22]. We run our method directly on masks predicted by RUP module without mask propagator (TMR). Overall, the dataset does not appear to be sufficiently challenging to differentiate between the methods.
@@ Appendix C Evaluation on RobustNeRF Dataset
Ours 25.10 0.84 22.90 0.87 34.25 0.95 33.85 0.93 32.45 0.93 29.71 0.90
Table 4: Quantitative comparison on the RobustNeRF dataset [21].
Anti-Stress Lab (1) Lab (2) Library Office Mean
WildGaussians w/o TMR 21.69 0.89 25.71 0.92 23.68 0.91 24.65 0.92 28.89 0.95 24.92 0.92
WildGaussians w/ TMR 24.07 0.92 24.65 0.92 24.84 0.93 28.32 0.97 29.75 0.96 26.33 0.94
Ours w/ TMR 28.79 0.97 27.71 0.95 25.42 0.93 28.34 0.97 29.87 0.96 28.03 0.96
Table 5: Evaluation of WildGaussians with TMR module on the Transient-3DGS dataset.
@@
Figure 6: Qualitative results on the On-the-go dataset using the training frames. Our method produces higher-quality renderings without artifacts.
@@
Figure 7: Qualitative results on the On-the-go dataset using the testing frames. Our method produces higher-quality renderings without artifacts.
Appendix D Additional Experiments with TMR Module
Even though our proposed TMR module leverages SAM2 to propagate the transient masks, we would like to emphasize that our method enables mask propagation spatially and temporally consistent, thereby providing more accurate and reliable reconstruction. Table 5 presents an evaluation of the reconstruction quality of WildGaussians with our TMR module. First, we obtained the transient masks using WildGaussians. Then, we propagate them through our TMR module. Finally, we reconstruct the scenes based on the transient masks obtained in the previous step. Our evaluation shows that our method produces higher-quality results for most scenes, with comparable performance in the remaining ones. Our method, with the TMR module, outperforms WildGaussians with the TMR module on Anti-Stress, Lab (1), Lab (2) scenes. The TMR module generally enhances the reconstruction quality of the original WildGaussians, but it is limited because of the false positive transient detections that come from WildGaussians itself. Furthermore, we note that the hyperparameters of our TMR module are highly dependent on the dataset rather than the model. That makes our TMR module robust across the different methods.
Appendix E More Qualitative Comparisons
@@
For qualitative comparison, we evaluate our method against SpotLessSplats [22], Robust3DGS [28], and WildGaussians [12]. We provide corresponding renderings for the masks shown in the main paper in Sec. 4.2. Fig. 6 and 7 show reconstructions of several scenes from the On-the-go dataset on training and testing frames, respectively. Currently, most methods can produce fairly good reconstructions and avoid significant artifacts, so generally, most methods produce fairly similar results (at least in the absence of semi-transient objects and other adversarial cases). Notably, compared to other residual-based methods, we avoid misclassifying high-frequency details and similar objects.
Figure 8: Comparison of predicted masks for full and reduced scenes.
@@
Figure 9: Comparison of predicted masks and scene reconstructions during the movement of semi-transient objects across different frames.
Appendix F Handling Semi-Transient Objects